OpenStudy (anonymous):

Check symmetry with respect to both axes and the origin of the graph of x^2y-x^2+4y=0

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

simple way is to replace x by -x and see if you get the same function then try replacing y by -y and see if you get the same function if thats the case it should be symm wrt to both axis or to either x if -x returns the same functin, or y if -y returns the same function

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

do you have the answers? lets verify the work

OpenStudy (anonymous):

Okay, thank you. I thought I should factor out the x^2 does that make sense?

OpenStudy (anonymous):

no, say you put x = -x then you will have (-x)^2y-(-x)^2+4y=0 which is same as original function as (-x)^2= x^2 so symettrical to x , then try for -y and it is not similar to original function

OpenStudy (anonymous):

Oh! okay, thank you. That makes sense, I thought maybe factoring would make the answer seem easier but that is what i was taught. Thank you so much!